Circular No. 8213

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2001 U 2 AND S/2002 N 4 A new Uranian satellite candidate (R = 24.9–25.1) was observed on

2001 Aug. 13 (Cerro Tololo, Blanco 4-m reflector) and Sept. 21 (Palomar, Hale 5-m refl.) by teams led by M. Holman and B. Gladman (for other participants and affiliations, see IAUC 7980), but further attempts to observe the object proved unsuccessful. On 2003 Sept. 24, B. G. Marsden identified this object, now designated S/2001 U 2, with a satellite candidate (R = 25.0–25.2) observed by S. S. Sheppard, University of Hawaii, on images obtained by D. C. Jewitt and himself on 2003 Aug. 29–30 (Subaru 8.2-m reflector) and Sept. 20 (Gemini 8.2-m reflector), and confirming observations were made by Holman on Sept. 30 (Las Campanas, Clay 6.5-m reflector).

A new Neptunian satellite candidate (R = 24.6–25.2) was observed on 2002 Aug. 14 and 16 (Blanco reflector) and Sept. 3 (Cerro Paranal, 8.2-m UT3 telescope) by teams led by Holman and Gladman (for other participants and affiliations see IAUC 8047) and also later on Sept. 3 (La Palma, 2.5-m Nordic Optical Telescope) by T. Grav. This object was tentatively recovered by Holman on 2003 Aug. 19 (Blanco reflector, R = 24.3) and confirmed by him on Sept. 28–30 (Clay reflector, R = 24.6).

Full astrometry, Marsden’s orbital elements (S/2001 U 2, a = 0.14 AU, e = 0.43, i = 167^{o}, H = 12.5; S/2002 N 4, a = 0.32 AU, e = 0.60, i = 139^{o}, H = 10.4), and ephemerides are on MPEC 2003-S105 and 2003-S107.

COMET C/2003 S4 (LINEAR) An apparently asteroidal object discovered by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been found to have cometary appearance by several CCD observers, including J. Young and A. Grigsby (Table Mountain, 0.6-m reflector; 5′′ round coma and faint 10′′ outer shell on Sept. 30.1–30.2 UT, possibly slightly extended in p.a. 145^{o} on Oct. 1.1–1.3); H. Mikuž (Crni Vrh, Slovenia; 0.60-m f/3 Deltagraph; diffuse coma of diameter ∼ 10′′ on Sept. 30.8–30.9; mag 18.1–18.2), and J. E. McGaha (Tucson, AZ, 0.62-m f/5.1 reflector; 2′′ coma elongated to the southeast on Oct. 1.2).

2003 UT  α_{2000}  δ_{2000}  Mag.
Sept.27.16008  21^{h}26^{m}06.^{s}41  +33^{o}07′10′′.6  19.3

The available astrometry, the preliminary parabolic orbital elements (T = 2004 June 6.4 TT, ω = 158^{o}.8, Ω = 223^{o}.9, i = 40^{o}.8, equinox 2000.0, q = 3.742 AU), and an ephemeris appear on MPEC 2003-T02.

2003 October 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green